OpenStudy (anonymous):

help me to write this out... the sum of three times a number, and 8, is equal to two times the number

OpenStudy (anonymous):

Okay the first part "the sum of three times a number" means the answer to 3 multiplied by a variable, let's call the variable, n. So this is what it would look like: \[3 \times n = \] Do you understand so far @diplomabound?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay cool! Now the next part "and 8" I'm guessing it's supposed to be in the same equation that I wrote above so this is what it would look like: \[3 \times 8n = \] And the next part, "is equal to two times a number" means that 2 is multiplied by a variable. Let's call this variable x. Now the equation looks like this: \[3 \times 8n = 2 \times n\]

OpenStudy (anonymous):

Oh sorry I forgot to add the x to that equation I'll change it: \[3 \times 8n = 2 \times x\]

OpenStudy (anonymous):

this is what I got 3(x+8)=2x

OpenStudy (anonymous):

your answer is different.

OpenStudy (anonymous):

hi welcome to Openstudy! :) "sum of three times a number and 8" should be \(\sf 3x + 8\) then the rest of your answers will be right ^_^

OpenStudy (anonymous):

Hi data_LG2 so then it should read 3x+8=2x? That's what I came up with

OpenStudy (anonymous):

yes, that would be right :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

one more question

OpenStudy (anonymous):

how would you write, the sum of the second and third even consecutive integers? is this correct:( x+2)+(x+4)